OpenStudy (calculusxy):

Draw a rectangle that is twice as wide as it is tall, and that fits snugly into the triangular region formed by the line 3x + 4y = 12 and the positive coordinate axes, with one corner at the origin and the opposite corner on the line. Find the dimensions of this rectangle.

OpenStudy (calculusxy):

@mathstudent55 @agent0smith

OpenStudy (calculusxy):

I usually have some difficulty understand when it says "twice as wide as it is tall," but I put the width to be x and the height to be 2x.

OpenStudy (calculusxy):

Is that correct?

OpenStudy (agent0smith):

"twice as wide as it is tall," means if the height is x, width is 2x

OpenStudy (calculusxy):

Okay. So 3x would be the width and the 4y would be the height right?

OpenStudy (calculusxy):

Like can I do 3(2x) + 4(x) = 12?

OpenStudy (agent0smith):

You should be using a diagram btw... The point on the line is (x, y), or (2x, x) Plug them into the equation 3x + 4y = 12 and solve for x, then find y.

OpenStudy (agent0smith):

Exactly.

OpenStudy (calculusxy):

6x + 4x = 12 10x = 12 x = 6/5

OpenStudy (calculusxy):

Width: 6/5 x 2 = 12/5 units Height: 6/5

OpenStudy (calculusxy):

Is that right?

OpenStudy (agent0smith):

Yep. A diagram would help confirm it though.

OpenStudy (calculusxy):

OK.

OpenStudy (calculusxy):

Can you help me with a couple more questions?

OpenStudy (agent0smith):

Okay

OpenStudy (calculusxy):

I made a new post and mentioned you too